Suppose we have the following data:

30

79

59

65

40

64

52

53

57

39

61

47

50

60

48

50

58

67

Suppose the number of nonoverlapping classes is determined to be 5. 

 

What is the mean and sample variance for this data?


1
Expert's answer
2021-05-12T03:52:25-0400

The distribution is-




Mean "\bar{x}=\dfrac{\sum xf}{\sum f}=\dfrac{970}{18}=53.88"


Sample Variance "=\dfrac{\sum (x-\bar{x})^2}{N-1}=\dfrac{1075.678}{18-1}=63.275"